OpenStudy (anonymous):

Using the fact that lim h->0 sin(h)/h =1 and lim h->0 cos(h)-1/h =0 compute the limit lim h->0 cos(x+h)-cos(x)/ h

OpenStudy (anonymous):

again?

OpenStudy (bradely):

lim h->0 cos(x+h)-cos(x)/ h =lim h->0 cosxcosh-sinxsinh-cos(x)/ h =lim h->0 cosx(cosh-1)-sinxsinh/ h =lim h->0( cosx(cosh-1)/h)-lim h->0sinxsinh/ h =0-sinx=-sinx Source: http://www.mathskey.com/question2answer/

OpenStudy (anonymous):

dang, guess it didn't work yesterday lets take it step by step \[\lim_{h\to 0}\frac{\cos(x+h)-\cos(x)}{h}=\lim_{x\to 0}\frac{\cos(x)\cos(h)-\sin(x)\sin(h)-\cos(x)}{h}\] is a start